The average words per minute a person can type is 50 words per minute with standard deviation of 2.5 words. If you choose 25 samples, what is the probability that the average words they can type is more than 39?
1
Expert's answer
2021-10-22T14:36:17-0400

Let "\bar{X}=" the average words they can type: "\bar{X}\sim N(\mu_{\bar{X}}, \sigma_{\bar{X}}^2)"

"\mu_{\bar{X}}=\mu=50, \sigma_{\bar{X}}=\sigma/\sqrt{n}=2.5/\sqrt{25}"


"P(\bar{X}>39)=1-P(\bar{X}\leq39)"

"=1-P(Z\leq\dfrac{39-50}{2.5/\sqrt{25}})=1-P(Z\leq-22)\approx1"